OpenStudy (anonymous):

8/9 - 12/9 = was wondering it -4/9 is it true?

OpenStudy (anonymous):

Look we have : 8/9-12/9=8-12/9=-4/9 So your answer is true !:)

OpenStudy (anonymous):

(8-12)/9 ie -4/9 yuup its true

OpenStudy (anonymous):

Thanks i wanted to be sure if i was correct or not/